# Asymptotic theory of in-context learning by linear attention

### v1

@@ 3.5 Memorization to ICL transition with increasing pretraining task diversity
$\lim_{\begin{subarray}{c}\tau\to\infty\\ \alpha\to\infty\end{subarray}}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}=\begin{cases}\rho+(1-\kappa)\left(1+\frac{\rho}{1+\rho}c^{\ast}\right)\qquad&\kappa<1\\ \rho&\kappa>1\end{cases}.$ (34)
This change in analytical behavior indicates a phase transition at $\kappa=1$. Further, the $\kappa>1$ branch approaches $\rho$, the error of the Bayes-optimal ridge estimator in this limit. The smooth memorization-ICL transition observed in Figure 3 for the finite $\alpha,\tau$ case stems from this phase transition.
4 Experiments with full, nonlinear Transformers
As our theoretical results are derived in a simplified setting, we aim to test if these insights are applicable to a full, nonlinear Transformer model. Specifically, we will investigate: (1) whether we have identified the correct scaling regime for non-trivial learning in an ICL task; (2) if the full Transformer exhibits a sample-wise double descent, and whether the location of the peak error scales quadratically with input dimensions as predicted by our theory; and (3) if the transition from memorization to generalization occurs, with the transition point being around $\kappa=1$.
(a) Nonlinear model exhibits double descent of test ICL error in scaling parameter $\tau$.
@@ 4 Experiments with full, nonlinear Transformers
(b) Interpolation threshold follows predicted $n\propto d^{2}$ scaling.
Figure 4: Experimental verification of both scaling definitions and double descent behaviour in $n$. Figure 4(a): Increasing $n$ will increase error until an interpolation threshold is reached.
@@
Figure 4(b) which occurs for $n$ proportional to $d^{2}$, as predicted by the linear theory. Best fit lines () correspond to fitting $\log(n)=a\log(d)+b$ giving $a_{2}=1.82,b_{2}=3.55$ for 2-layer model and $a_{3}=2.22,b_{3}=2.81$ for 3-layer model. Interpolation threshold was computed empirically by searching for location in $\tau$ of sharp increase in value and variance of training error at a fixed number of gradient steps. Parameters: $d=10,\alpha=5,\kappa=\infty,\rho=0.25^{2}$. For fig. 4(a): 2-layer architecture; variance shown comes from model trained over different samples of pretraining data; lines show averages over 10 runs and shaded region shows standard deviation.
@@
Our experiments [^1] Code to reproduce all experiments is available at https://github.com/Pehlevan-Group/icl-asymptotic. are done with a standard Transformer architecture consisting of blocks with: (1) a single-head softmax self-attention with $K,Q,V\in\mathbb{R}^{hd\times\ell(d+1)}$ matrices [^2] Provided that the hidden layer dimension $hd$ is greater than $d$, it does not affect the expressivity of the attention mechanism. In our experiments, we use $h=10$., followed by (2) a two-layer dense MLP with GELU activation and hidden layer of size $10d$ [1]. Residual connections are used between the input vector, the pre-MLP output, and the MLP output. Each sample takes the form given by eq. 4. We use either two or three Transformer blocks before returning the final logit corresponding to the $(d+1,\ell+1)$th element in the embedding. The loss function is the mean squared error (MSE) between the predicted label (the output of the model for a given sample $Z$) and the true $\hat{y}_{\ell+1}$ value. We train the model in an offline setting with $n$ samples $Z_{1},\cdots,Z_{n}$, using the Adam optimizer [40] with a learning rate $10^{-4}$ until the training error converges, typically requiring 300000-500000 gradient steps. The structure of the pretraining and test distributions exactly follows the setup described in Section 2.
In Figure 4a, we plot the generalization error and observe the double descent behavior of a full, nonlinear Transformer for the ICL task as the number of pretraining samples varies (plotted as a function of $\tau$). We find that the peak of this curve occurs at the interpolation threshold, identified by tracking when the training loss is non-zero (see figure caption). Our theory predicts that the number of samples $n$ at the peak of the curve, as well as the interpolation threshold, should scale with $d^{2}$. This scaling is indeed observed in Figure 4b for the full, nonlinear Transformers. These observations suggest that the nonlinear Transformer operates within the scaling regime we have identified.
Figure 5: Experiment in 2-layer nonlinear network demonstrates a sharp transition between dMMSE estimator and ridge estimator, familiar from the linear theory. Parameters: $d=20,\tau=50,\alpha=5,\rho=0.01$. Variance shown comes from model trained over different samples of pretraining data; lines show averages over 10 runs and shaded region shows standard deviation.
Further, as observed before in Raventós et al. [18], we recover the memorization-to-ICL transition as a function of pretraining task diversity, shown in Figure 3. We note that this transition happens near $\kappa=1$ in the nonlinear model, consistent with our theoretical predictions on the linear model.
5 Discussion
@@
In this paper, we computed sharp asymptotics for the in-context learning (ICL) performance in a simplified model of ICL for linear regression using linear attention. This exactly solvable model demonstrates a transition from memorizing pretraining task vectors to generalizing beyond them as the diversity of pretraining tasks increases, echoing empirical findings in full Transformers [18]. Additionally, we observe a sample-wise double descent as the amount of pretraining data increases. Our numerical experiments show that full, nonlinear Transformers exhibit similar behavior in the scaling regime relevant to our solvable model. Our work represents a first step towards a detailed theoretical understanding of the conditions required for ICL to emerge [19].
@@
Our paper falls within a broader program of research that seeks sharp asymptotic characterizations of the performance of machine learning algorithms. This program has a long history in statistical physics [41, 42, 38], and has in recent years attracted substantial attention in machine learning [36, 43, 44, 45, 46, 47, 48, 38, 49]. For simplicity, we have assumed that the covariates in the in-context regression problem are drawn from an isotropic Gaussian. However, our technical approach could be extended to anisotropic covariates, and, perhaps more interestingly, to featurized linear attention models in which the inputs are passed through some feature map before linear attention is applied [33, 32]. This extension would be possible thanks to an appropriate form of Gaussian universality: for certain classes of regression problems, the asymptotic error coincides with that of a model where the true features are replaced with Gaussian features of matched mean and covariance [36, 43, 44, 45, 46, 47, 48, 37, 50]. This would allow for a theoretical characterization of ICL for realistic data structure in a close approximation of full softmax attention, yielding more precise predictions of how performance scales in real Transformers.
@@
Finally, our results have some bearing on the broad question of what architectural features are required for ICL [7, 11, 19]. Our work shows that a full Transformer—or indeed even full linear attention—is not required for ICL of linear regression. However, our simplified model retains the structured quadratic interaction between inputs that is at the heart of the attention mechanism. It is this quadratic interaction that allows the model to solve the ICL regression task, which it does essentially by reversing the data correlation. One would therefore hypothesize that our model is minimal in the sense that further simplifications within this model class would impair its ability to solve this ICL task. In the specific context of regression with isotropic data, a simple point of comparison would be to fix $\Gamma=I_{d}$, which gives a pretraining-free model that should perform well when the context length is very long. However, this further-reduced model would perform poorly if the covariates of the in-context task are anisotropic. More generally, it would be interesting to investigate when models lacking this precisely-engineered quadratic interaction can learn linear regression in-context, and if they are less sample-efficient than the attention-based models considered here.
Acknowledgements
YML was supported by NSF Award CCF-1910410, and by the Harvard FAS Dean’s Fund for Promising Scholarship. JAZV and CP were supported by NSF Award DMS-2134157 and NSF CAREER Award IIS-2239780. CP is further supported by a Sloan Research Fellowship. AM acknowledges support from Perimeter Institute, which is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. This research was supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), through the authors’ participation in the Fall 2023 program “Deep Learning from the Perspective of Physics and Neuroscience.”

### v2

@@ III.4 ICL and IDG error curves can have non-monotonic dependence on context length
$\lim_{\alpha\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ =$=\begin{dcases}\infty&\kappa\leq\text{min}(\tau,1),\\ 1-\tau+\rho+\frac{\rho\kappa\tau}{(\kappa-\tau)(1-\tau)}&\tau<1,\kappa>\tau,\\ \rho+\frac{\rho\kappa}{(\kappa-1)(\tau-1)}&\tau>1,\kappa>1.\end{dcases}$ (35)
Note the explicit divergence of ICL error for $\kappa\leq\text{min}(\tau,1)$ (independent of $\rho$). Similar investigation of Result 2 under the limit $\alpha\to\infty$ reveals an explicit divergence of IDG error for $\kappa=\tau$. The divergence in $\alpha$ of ICL error is a peculiarity of the ridgeless limit of errors we have taken in Results 1 and 2, and is further discussed in Section SI.5.4 of the Supplementary Information.
III.5 Learning transition with increasing pretraining task diversity
((a)) $g_{\text{task}}$ for various $\alpha$, log-log scale
((b)) $g_{\text{task}}$ for various $\alpha$, log-linear scale
Figure 4: Plots of $g_{\text{task}}$ against $\kappa$ for linear transformer model vs dMMSE estimator. Figure (a) has plots on log-log scale highlighting the substantial difference in their rates of convergence towards 0. Dashed and solid lines are theory predictions; dots and squares are numerical simulation. Figure (b) has plots on log-linear scale with the addition of the green $\alpha\to\infty$ curve given by (36) demonstrating the phase transition at $\kappa=1$. Parameters: Linear transformer: $d=100-140$, $\tau=0.2\alpha$ throughout, simulations computed over 20 runs; dMMSE: $d=100$, simulations computed over 1000-5000 runs
It is important to quantify if and when a given model is actually learning in-context, i.e., solving a new regression problem by adapting to the specific structure of the task rather than relying solely on memorized training task vectors. We refer to this phenomenon as task generalization.
We expect pretraining task diversity ($k$) to play a key role in task generalization. To provide some motivation for this, suppose the model was pretrained on only a single task vector ($k=1$). It is reasonable to expect that the model will memorize this task vector and make predictions under the prior that the task vector in the test examples is identical to that particular pretraining task vector. On the other extreme, suppose the model was pretrained by drawing a fresh task vector for every training sequence. Here, the expectation would be that the model achieves task generalization. Indeed, results from previous computational studies on the linear regression task [20] suggest that transformers are better in-context learners if they have seen a greater task diversity in the training set, and further the models transition from a task memorization phase to a task generalization phase as task diversity increases. Hence, we study the task generalization behavior of the linear transformer as the task diversity increases and precisely quantify the task diversity scale at which such transition happens.
First, we need to discuss how to measure task generalization. To motivate our solution, we take a closer look at the ridge estimator ($w_{\text{Bayes}}^{\text{ridge}}$ as in eq. 24), which is the Bayes-optimal estimator given the ICL task structure. We analytically characterize $e_{\text{ICL}}$ and $e_{\text{IDG}}$ errors for this estimator in SI Section SI.6, giving
$e^{\text{ridge}}_{\text{ICL}}=\rho\left(1+\frac{1}{\alpha}\mathcal{M}_{\alpha}\left(\frac{\rho}{\alpha}\right)\right)=e^{\text{ridge}}_{\text{IDG}}.$
Note that the performance of the ridge estimator is identical on the ICL and IDG test distributions, meaning that it does not rely at all on memorizing task vectors.
Motivated by these arguments, we propose to measure the task generalization capability of a model by studying the difference between its performance on the ICL test distribution and its performance on the IDG test distribution. Specifically, we consider the quantity $g_{\text{task}}=e_{\text{ICL}}-e_{\text{IDG}}$ for a given model or estimator. This difference being large implies the model, performing better on training tasks, has not learned the true task distribution and is not generalizing in task. Conversely, a small difference between ICL error and IDG error suggests that the model is leveraging the underlying structure of the task distribution rather than overfitting to and interpolating specific task instances seen in training.
Simulations of $g_{\text{task}}$ as a function of task diversity, $\kappa=k/d$, are shown in Figure 4 for two inference models: dMMSE estimator, and the linear transformer. The dMMSE estimator is included as a benchmark quantifying the performance of a “pure memorizer”, as it is the Bayes-optimal estimator that assumes task vectors only seen in the training test. We see that, as the task diversity parameter $\kappa$ increases, $g_{\text{task}}$ falls for both estimators but at very different rates.
The $g_{\text{task}}$ for the dMMSE estimator falls very slowly as a function of $\kappa$. This is expected because the explicit form of $w_{\text{Bayes}}^{\text{dMMSE}}$, given in eq. 25, is that of a kernel smoother employing a Gaussian kernel as the weighting function around each task vector. Known results about this class of estimators show the performance of the dMMSE estimator on the ICL task suffers from the “curse of dimensionality” [42, 43, 44], i.e. the required number of training samples $k=d\kappa$ would need to be exponential in $d$ to outperform some given $e_{\text{ICL}}$ error tolerance. Intuitively, the IDG task distribution $\mathcal{P}_{\text{test}}=\mathcal{P}_{\text{train}}=\text{Unif}\{w_{1},\cdots,w_{k}\},w_{i}\sim\mathcal{N}(0,I_{d})$ and the ICL task distribution $\mathcal{P}_{\text{test}}=\mathcal{N}(0,I_{d})$ are indistinguishable (by appropriate measures) only if $k$ is exponentially large in $d$. We present more formal arguments in SI Section SI.6.
On the other hand, the $g_{\text{task}}$ for the linear transformer estimator first behaves similarly to the dMMSE estimator, but starts falling sharply after around $\kappa=1$. How quickly does $g^{\text{LT}}_{\text{task}}$ limit to 0 as $\kappa\to\infty$? By expanding $e_{\text{ICL}}-e_{\text{IDG}}$ from (31) and (32) in $\kappa$ we have the leading asymptotic behavior given by
$g^{\text{LT}}_{\text{task}}=0+\frac{1}{\kappa}\begin{cases}c_{1}&\tau<1\\ c_{2}&\tau>1\\ \end{cases}+\mathcal{O}\left(\frac{1}{\kappa^{2}}\right)$
for constants $c_{1},c_{2}$ dependent on $\alpha,\tau,\rho$, provided in SI Section SI.6. This sharp fall signals a transition of the model to the task generalization regime. To further understand the nature of this transition and the role of $\kappa$ in the solution learned by the linear attention mechanism, consider the regime where $\tau,\alpha\to\infty$ with $\tau/\alpha\equiv c^{\ast}$ kept fixed. Under this setting, we have
$\lim_{\begin{subarray}{c}\tau\to\infty\\ \alpha\to\infty\end{subarray}}g^{\text{LT}}_{\text{task}}=\begin{cases}(1-\kappa)\left(1+\frac{\rho}{1+\rho}c^{\ast}\right)\qquad&\kappa<1\\ 0&\kappa>1\end{cases}\,.$ (36)
This change in analytical behavior indicates a phase transition at $\kappa=1$. We see this phenomena in both Figure 4(a) and Figure 4(b) where the $g_{\text{task}}$ curve plotted against $\kappa$ becomes nondifferentiable at $\kappa=1$ for $\alpha\to\infty$.
We conclude that the linear transformer model is a much more efficient task generalizer than the dMMSE estimator which is the Bayes-optimal memorizer, as shown in Figure 4. The $1/\kappa$ decay in $g^{\text{LT}}_{\text{task}}$ vs the much slower decay of $g^{\text{dMMSE}}_{\text{task}}$ suggests that the linear transformer quickly learns an inference algorithm that generalizes in-context rather than interpolates between training tasks.
IV Experiments with fully-parameterized nonlinear Transformers
((a)) Error curve in $\tau$: Linear attention, full $K,Q,V$ matrices
((b)) Error curve in $\tau$: Nonlinear model, softmax attention only
((c)) Error curve in $\tau$: Nonlinear model, two blocks of softmax attention and 1-layer MLP
@@ IV Experiments with fully-parameterized nonlinear Transformers
((d)) Interpolation threshold follows predicted quadratic $n\propto d^{2}$ scaling in a range of linear and nonlinear architectures.
@@
Figure 5: Experimental verification in full linear attention (a) and nonlinear models (b), (c), of both scaling definition for $\tau$ and double descent behavior in $n$. 5(a), 5(b), 5(c) show error curves against $\tau$ for various architectures, consistent across token dimension $d=20,40,80$. Some deviations arise near double descent in the linear model (a), possibly due to parameter inversion or instabilities, and for large $\tau$ in the deep nonlinear model (c), possibly due to the training procedure. Double-descent phenomena are confirmed: increasing $n$ will increase error until an interpolation threshold is reached. colored dashed lines indicate experimental interpolation threshold for that architecture and $d$ configuration. 5(d) shows that the location of the interpolation threshold occurs for $n$ proportional to $d^{2}$, as predicted by the linear theory. Dots are experimental interpolation thresholds for various architectures, and dashed lines are best fit curves correspond to fitting $\log(n)=a\log(d)+b$, each with $a\approx 2$ (explicitly, $a_{\text{full linear}}=1.87$, $a_{\text{softmax}}=1.66$, $a_{\text{2 blocks}}=2.13$, $a_{\text{3 blocks}}=2.08$). Interpolation threshold was computed empirically by searching for location in $\tau$ of sharp increase in value and variance of training error at a fixed number of gradient steps. Parameters: $\alpha=1,\kappa=\infty,\rho=0.01$. For Figure 5(a), 5(b), and 5(c): variance shown comes from model trained over different samples of pretraining data; lines show averages over 10 runs and shaded region shows standard deviation.
Given that our theoretical results are derived from a simplified linear attention setting, we aim to determine if these insights can be extended to a nonlinear and deep Transformer models with full $K,Q,V$ parameterization. Even though we would not expect the specific algorithm discussed in Section II.4 to transfer to these nonlinear settings, we will test whether our proposed scalings and main qualitative predictions remain. Specifically, we will investigate the following: (1) whether we have identified the correct scaling regime for non-trivial learning in an in-context learning (ICL) task; (2) whether the full Transformer exhibits a sample-wise double descent, where the location of the peak error scales quadratically with input dimensions as predicted by our theory; and (3) whether the transition from memorization to task generalization occurs.
Figure 6: Plot of $g_{\text{task}}$ against $\kappa$ (loglinear scale) for the nonlinear architecture given in Figure 5(c). This demonstrates consistency of $\kappa$ scaling across increasing dimension choices $d=20,40,80$, as well as a transition in task-generalization measure $g_{\text{task}}$ near $\kappa=1$, familiar from the linear theory. Parameters: $\tau=10,\alpha=1,\rho=0.01$. Variance shown comes from 10 models trained over different samples of pretraining data.
@@
Our experiments [^1] Code to reproduce all experiments is available at https://github.com/Pehlevan-Group/incontext-asymptotics-experiments. are done with a standard Transformer architecture, where each sample context initially takes the form given by (4). The fully-parameterized linear transformer (Figure 5(a)) and the softmax-only transformer (Figure 5(b)) do not use a multilayer perception (MLP) in addition to attention. If MLPs are used (e.g. Figure 5(c) and Figure 5(d)), the architecture consists of blocks with: (1) a single-head softmax self-attention with $K,Q,V\in\mathbb{R}^{(d+1)\times(d+1)}$ matrices, followed by (2) a two-layer dense MLP with GELU activation and hidden layer of size $d+1$ [1]. Residual connections are used between the input tokens (padded from dimension $d$ to $d+1$), the pre-MLP output, and the MLP output. We use a variable number of attention+MLP blocks before returning the final logit corresponding to the $(d+1,\ell+1)$th element in the original embedding structure given by (4). The loss function is the mean squared error (MSE) between the predicted label (the output of the model for a given sample $Z$) and the true value $y_{\ell+1}$. We train the model in an offline setting with $n$ total samples $Z_{1},\cdots,Z_{n}$, divided into 10 batches, using the Adam optimizer [45] with a learning rate $10^{-4}$ until the training error converges, typically requiring 10000 epochs [^2] Note that larger $d$ models are often trained for fewer epochs than smaller $d$ models due to early stopping; that said, whether or not early stopping is used in training does not affect either the alignment of error curves in $d$-scaling nor the qualitative behavior (double descent in $\tau$ and transition in $\kappa$.. The structure of the pretraining and test distributions exactly follows the setup for the ICL and IDG tasks described in Section II.
In Figure 5, we plot the average error for a range of fully-parameterized linear (Figure 5(a)) and nonlinear architectures (Figure 5(b) and Figure 5(c)) on the ICL task against the parameter $\tau=n/d^{2}$, which measures the pretraining sample complexity. Notice that the large-$d$ full $K,Q,V$ parameterized linear architecture studied in Figure 5(a) follows the general trend of the theory derived on a simplified parameterization. Further, in all figures, plots across different values of $d$ roughly collapse, suggesting again that our scaling relations capture the observed behavior.
Across all architectures, we observe a double descent in $\tau$. The peak of this curve occurs at the interpolation threshold, which we identify by tracking the value of $\tau$ at which the training loss becomes non-zero (see figure caption). Our theory predicts that the number of samples $n$ at the peak, as well as the interpolation threshold, should scale with $d^{2}$. Indeed, for the fully-parameterized linear transformer in Figure 5(a), the double descent occurs very close to $\tau=1$ for larger $d$, as predicted by the reduced-parameter linear theory. This general quadratic scaling is summarized in Figure 5(d) across a range of architectures. These findings suggest that nonlinear Transformers exhibit a similar interpolation transition, with the same quadratic scaling in sample size as seen in the simplified model.
Finally, we can investigate the effects of task diversity on $g_{\text{task}}$ in a nonlinear model, shown in Figure 6. This provides empirical evidence towards the choice of $\kappa$ scaling, as errors are consistent over a range of token dimension choices. Further, we recover the memorization-to-task-generalization transition in $\kappa$, as observed before in [20]. In Figure 6, we see the value of $\kappa$ for which $g_{\text{task}}$ appears to converge towards 0 is below 1 (as in (36)), suggesting that the more powerful architecture used here may be more “task-efficient” than the linear architecture in achieving memorization-to-generalization transition.
V Discussion
@@
In this paper, we computed sharp asymptotics for the in-context learning (ICL) performance in a simplified model of ICL for linear regression using linear attention. This exactly solvable model demonstrates a transition from memorizing pretraining task vectors to generalizing beyond them as the diversity of pretraining tasks increases, echoing empirical findings in full Transformers [20]. Additionally, we observe a sample-wise double descent as the amount of pretraining data increases. Our numerical experiments show that full, nonlinear Transformers exhibit qualitatively similar behavior in the scaling regime relevant to our solvable model. Understanding the mechanistic underpinnings of ICL of well-controlled synthetic tasks in solvable models is an important prerequisite to understanding how it emerges from pretraining on natural data [21].
@@
Our paper falls within a broader program of research that seeks sharp asymptotic characterizations of the performance of machine learning algorithms. This program has a long history in statistical physics [46, 47, 41], and has in recent years attracted substantial attention in machine learning [38, 48, 49, 50, 51, 52, 53, 41, 54]. For simplicity, we have assumed that the covariates in the in-context regression problem are drawn from an isotropic Gaussian distribution. However, our technical approach could be extended to anisotropic covariates, and, perhaps more interestingly, to featurized linear attention models in which the inputs are passed through some feature map before linear attention is applied [35, 34]. This extension would be possible thanks to an appropriate form of Gaussian universality: for certain classes of regression problems, the asymptotic error coincides with that of a model where the true features are replaced with Gaussian features of matched mean and covariance [38, 48, 49, 50, 51, 52, 53, 39, 55]. This would allow for a theoretical characterization of ICL for realistic data structure in a closer approximation of full softmax attention, yielding more precise predictions of how performance scales in real Transformers.
@@
Finally, our results have some bearing on the broad question of what architectural features are required for ICL [7, 13, 21]. Our work shows that a full Transformer—or indeed even full linear attention—is not required for ICL of linear regression. However, our simplified model retains the structured quadratic pairwise interaction between inputs that is at the heart of the attention mechanism. It is this quadratic interaction that allows the model to solve the ICL regression task, which it does essentially by reversing the data correlation (see §II.4). One would therefore hypothesize that our model is minimal in that further simplifications within this model class would impair its ability to solve this ICL task. In the specific context of regression with isotropic data, a simple point of comparison would be to fix $\Gamma=I_{d}$, which gives a pretraining-free model that should perform well when the context length is very long. However, this further-reduced model would perform poorly if the covariates of the in-context task are anisotropic. Generally, it would be interesting to investigate when models lacking this precisely-engineered quadratic interaction can learn linear regression in-context [57], and if they’re less sample-efficient than the attention-based models considered here.
Acknowledgments
Funding acknowledgments
We thank William L. Tong for helpful discussions regarding numerics. YML was supported by NSF Award CCF-1910410, by the Harvard FAS Dean’s Fund for Promising Scholarship, and by a Harvard College Professorship. JAZV and CP were supported by NSF Award DMS-2134157 and NSF CAREER Award IIS-2239780. JAZV is presently supported by a Junior Fellowship from the Harvard Society of Fellows. CP is further supported by a Sloan Research Fellowship. AM acknowledges support from Perimeter Institute, which is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. This research was supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), through the authors’ participation in the Fall 2023 program “Deep Learning from the Perspective of Physics and Neuroscience.”
Author contributions
Y.M.L. led formal analysis, along with M.L., J.A.Z-V., A.M., and C.P.   M.L. led computational results and verification. C.P. and Y.M.L supervized the project. C.P. conceptualized the project. All authors contributed to key discussions and manuscript drafting.